# Iso-length-spectral Hyperbolic Surface Amalgams

### v1

@@ Abstract
Two negatively curved metric spaces are iso-length-spectral if they have the same multisets of lengths of closed geodesics. A well-known paper by Sunada provides a systematic way of constructing iso-length-spectral surfaces that are not isometric. In this paper, we construct examples of iso-length-spectral surface amalgams that are not isometric, generalizing Buser’s combinatorial construction of Sunada’s surfaces. We find both homeormophic and non-homeomorphic pairs. Finally, we construct a noncommensurable pair with the same weak length spectrum, the length set without multiplicity.
@@ 1 Introduction
Recall that the (unmarked) length spectrum of a metric space $(X,g)$ is an ascending multiset of positive real numbers representing lengths of closed geodesics in $(X,g)$. We say that two metric spaces are iso-length-spectral if they have the same length spectra.
The question of whether there exist iso-length-spectral manifolds that are not isometric has a long history. The first examples of such objects are due to Milnor (see [Mil64]), who constructed two flat, 16-dimensional tori. In [Vig80], Vignerás constructed examples of hyperbolic surfaces that are iso-length-spectral but not isometric. Both Vignerás’s and Milnor’s constructions are heavily number theoretic and, one could argue, difficult to describe geometrically. In [Sun85], Sunada developed a celebrated, systematic, and more geometric way of constructing hyperbolic iso-length-spectral surfaces. In [Bus86] (and [Bus92] for the genus $6$ case), Buser uses [Sun85] to construct pairs of iso-length-spectral, non-isometric surfaces of genus $\geq 5$.
@@
We remark that in contrast to the surface amalgam case, there cannot exist non-homeomorphic iso-length-spectral pairs of hyperbolic closed surfaces, as the length spectrum completely determines the genus of a surface (see [Cha84]). In the setting of $3$-manifolds, on the other hand, an applications of the Sunada method (e.g. [Spa89], [Rei92], [McR14]) have been shown to yield hyperbolic, iso-length-spectral 3-manifolds which are non-isometric and thus non-homeomorphic due to Mostow Rigidity. This provides further evidence towards the fact that surface amalgams share characteristics of both surfaces and 3-manifolds (see also [HST20]).
@@
We follow the terminology from [PR15] and define the weak length spectrum of a locally CAT($-1$) metric space to be a collection of lengths of closed geodesics without multiplicity. If two metric spaces have the same weak length spectrum, we say they are weak length isospectral. Notice that since the weak length spectrum is a subset of the length spectrum, weak length isospectrality is a weaker condition than length isospectrality.
@@
Acknowledgements. I am deeply indebted to Chris Leininger for suggesting the project, discussing it in depth with me, and pointing out gaps in previous drafts. I also thank Alan Reid for discussions that inspired Theorem 1.3, and for pointing out various helpful references. Finally, I thank my doctoral advisors Tullia Dymarz and Caglar Uyanik for helpful discussions and moral support. This work was partially supported by an NSF RTG grant DMS-2230900.
@@ 2 Buser’s Techniques
We rely heavily on the techniques from [Bus86], which we briefly sketch. Buser constructs two iso-length-spectral but non-isometric hyperbolic genus 5 surfaces by gluing together identical copies of right-angled octagons, which he calls building blocks. The gluing scheme is shown below in Figure 1. His construction is modeled off Sunada’s construction using almost conjugate subgroups of $(\mathbb{Z}/8\mathbb{Z})^{\times}\ltimes(\mathbb{Z}/8\mathbb{Z})^{+}$ given by Gerst in [Ger70] (see Example 1 of Section 1 of [Sun85]).
@@
Let $10<c<b<$ 1. Then any geodesic curve $\delta$ of $B$ which connects two sides of $B$ has length $\ell(\delta)\geq c$. Equality holds only if $\delta$ is a side (of length $c$) of a building block.
@@
Buser’s surfaces are iso-length-spectral. Next, for every closed geodesic in $S_{1}$, Buser constructs one in $S_{2}$ with the same length (and vice versa) using a technique he calls transplantation. Given a closed geodesic $\gamma\subset S_{1}$ with a starting point $p$ in the interior of some building block $B_{i}\subset S_{1}$, Buser specifies which building block $B^{\prime}_{i}\subset S_{2}$ one should start constructing $\gamma^{\prime}\subset S_{2}$ in so that $\ell(\gamma)=\ell(\gamma^{\prime})$. He uses the following set of rules, which depend on the parities of #$\#a$ and #$\#b$, defined below:
@@
Let #$\#a$ and #$\#b$ be the number of times a curve $\gamma$ transversely crosses sides of lengths $a$ and $b$ respectively, and let $B_{n_{1}}$ and $B^{\prime}_{n^{\prime}_{1}}$ denote the building blocks in $S_{1}$ and $S_{2}$ respectively in which to initiate $\gamma^{\prime}$.
@@
    If #$\#a$ is even, take $n_{1}=n^{\prime}_{1}$;
@@
    If #$\#a$ is odd and #$\#b$ is even, take $n^{\prime}_{1}=n_{1}+1$;
@@
    If #$\#a$ and #$\#b$ are both odd, take $n^{\prime}_{1}=n_{1}+2$.
@@
To construct $\gamma^{\prime}$, one simply decomposes $\gamma$ into geodesic segments $\{\gamma_{j}\}_{i=1}^{N}$ such that each $\gamma_{j}$ is contained completely in some building block $B_{j}$ and $\gamma_{1}$ and $\gamma_{N}$ are contained in the same building block and share an endpoint at $p$. Then, for each $\gamma_{j}$, taking advantage of the fact that the building blocks in $S_{1}$ and $S_{2}$ are all identical, one can construct a $\gamma^{\prime}_{j}\subset B^{\prime}_{j}$ identical to each $\gamma_{j}\subset B_{j}$. Buser shows that following Algorithm 2.2, the set of $\{\gamma^{\prime}_{j}\}$ will always close up to a geodesic loop. Furthermore, since $\ell(\gamma_{j})=\ell(\gamma^{\prime}_{j})$ for all $j$, it follows that $\ell(\gamma)=\ell(\gamma^{\prime})$, as desired. The same set of initiation rules also applies in the other direction (constructing a geodesic in $S_{1}$ given one in $S_{2}$).
@@ 3.1 Homeomorphic, non-isometric surface amalgams
Consider $S_{1}$ and $S_{2}$ from [Bus86]. In each surface, consider the four closed geodesics of length $2a$ obtained from concatenating the top right edge of length $a$ on $B_{2}$ (resp. $B_{6}$) with the top left edge of length $a$ on $B_{3}$ (resp. $B_{7}$). We do the same for the bottom edges. We then identify all the red edges of length $a$ on $B_{2}$ and $B_{6}$ (resp. $B^{\prime}_{2}$ and $B^{\prime}_{6}$) from Figure 2 and all blue edges on $B_{3}$ and $B_{7}$ (resp. $B^{\prime}_{3}$ and $B^{\prime}_{7}$).
Figure 2: Homeomorphic, non-isometric, iso-length-spectral surface amalgams. The gluing curve consists of four closed geodesics of length $2a$ identified together. Geodesic segments that are the same colors in $X_{1}$ are identified; the same is true for $X_{2}$. Cutting along the gluing geodesics and the systoles yields two connected components (in yellow and blue) for $X_{1}$ but only one connected component for $X_{2}$.
@@
$X_{1}$ and $X_{2}$ are iso-length-spectral. Consider $\gamma$, a closed geodesic in $X_{1}$. Decompose $\gamma$ into $\bigcup\limits_{i=1}^{N}\beta_{i}$ so that each $\beta_{i}$ is a continuous geodesic segment in the original surface $S_{1}$ with endpoints on the gluing curve. Furthermore, we require that none of the interiors of the $\beta_{i}$’s intersect the gluing curve. We say a closed geodesic in $\gamma_{1}\subset X_{1}$ is identical to $\gamma$ if $\gamma_{1}$ and $\gamma$ begin at the same point on the gluing curve $x\in X_{1}$, and $\gamma_{1}$ can be decomposed into $\bigcup\limits_{i=1}^{N}\beta_{i}^{1}$, a union of translated copies of $\beta_{i}$ in $S_{1}$ with endpoints on the gluing curve. There is a natural isometry $\varphi$ between gluing geodesics in $X_{1}$ and those in $X_{2}$ (here, we have the identity isometry). We say $\gamma^{\prime}\subset X_{2}$ is identical to $\gamma\subset X_{1}$ if $\gamma^{\prime}$ begins and ends at $\varphi(x)\in X_{2}$ while $\gamma$ begins and ends at $x\in X_{1}$, and $\gamma_{2}$ can be decomposed into $\bigcup\limits_{i=1}^{N}\beta^{\prime}_{i}$ so that each $\beta^{\prime}_{i}$ is a transplanted copy of $\beta_{i}$ (in the sense of [Bus86]) that begins and ends on the gluing curve in $X_{2}$. We will argue that the numbers of closed geodesics identical to $\gamma$ in $X_{1}$ and $X_{2}$ are the same, which shows the length spectra are the same.
@@
Figure 4: Two examples of (parts of) inadmissible $\beta_{i}$’s. In (a), $\beta_{i-1}\cup\beta_{i}$ is a backtracking geodesic. In (b), $\beta_{i-1}$ and $\beta_{i}$ begin at different points of the gluing curve, as the red and blue segments are not identified, which means $\beta_{i-1}\cup\beta_{i}$ is disconnected.
We claim that in $X_{1}$, there is one translate of $\beta_{i}^{1}:=\beta_{i}$, $\beta_{i}^{2}$, which begins and ends at the same point on the gluing curve $\beta_{i}$ begins and ends at. If $\beta_{i}$ begins on the top right (resp. top left, bottom right, or bottom left) side of length $a$ in $B_{i}$, initiate the second translate on the top right (resp. top left, bottom right, or bottom left) side of length $a$ in $B_{i+4}$. Let $\beta_{i}^{j}=\bigcup\limits_{k=0}^{L}\gamma_{i,k}^{j}$ be a decomposition of $\beta_{i}^{j}$ into geodesic segments that are contained entirely in a single building block, where $\gamma_{i,k}^{j}\subset B_{n_{i,k}^{j}}$.
@@
$c_{i}$ :=#$:=\#\{\beta^{j}_{i}\text{ which can be concatenated with some fixed copy of }\beta_{i-1}\}$
@@
We now examine how to choose a copy of $\beta_{N}$ from $\{\beta_{N}^{j}\}_{j=1}^{2}$. The set of admissible $\beta_{N}$ now depend on our choice of $\beta_{N-1}^{j}$ and $\beta_{1}^{1}$, and specifically on whether the following facts are true:
  • Fact 1: $\beta_{N-1}$ ends on a top red edge and $\beta_{N}$ begins on a top red edge (or the statement is true if we replace “red” with “blue” and/or “top” with “bottom”);
@@
Case One: Neither Fact 1 nor Fact 2 is true. In this case, any choice of pairs of translates of $\beta_{1}$ and $\beta_{N}$ is compatible. Thus, each choice of $\beta_{1}^{j}$ ($j=1,2$) has $\prod\limits_{i=2}^{N}c_{i}=\prod\limits_{i=2}^{N-1}(2)$ choices of sequences of admissible translates of $\beta_{i}$. Then $C_{1}(\gamma)=2\bigg{(}\prod\limits_{i=2}^{N-1}c_{i}\bigg{)}(2)$. See Figure 5.
Figure 5: An illustration of Case 1, with identical copies of closed geodesics created by concatenating the pink, orange, and green geodesic segments. Geodesic segments beginning on the top (resp. bottom) building blocks are denoted by solid (resp. dashed) lines. Here, $N=3$ and $c_{2}=2$, so $C_{1}(\gamma)=2(2)(2)=C_{2}(\gamma)$. We can check that there are indeed eight identical, non-backtracking closed geodesics in each of $X_{1}$ and $X_{2}$.
@@
Figure 6: An illustration for Case 2, with the same closed geodesic as before except with modified pink and dark green segments. Here, $N=3$ and $c_{2}=2$, so $C_{1}(\gamma)=2(2)=C_{2}(\gamma)$. We can check that there are indeed four identical, non-backtracking closed geodesics in each of $X_{1}$ and $X_{2}$.

### v2

@@ Abstract
Two negatively curved metric spaces are iso-length-spectral if they have the same multisets of lengths of closed geodesics. A well-known paper by Sunada provides a systematic way of constructing iso-length-spectral surfaces that are not isometric. In this paper, we construct examples of iso-length-spectral surface amalgams that are not isometric, generalizing Buser’s combinatorial construction of Sunada’s surfaces. We find both homeomorphic and non-homeomorphic pairs. Finally, we construct a noncommensurable pair with the same weak length spectrum, the length set without multiplicity.
@@ 1 Introduction
Recall that the (unmarked) length spectrum of a metric space $(X,g)$ is the ascending multiset of positive real numbers representing lengths of closed geodesics in $(X,g)$. We say that two metric spaces are iso-length-spectral if they have the same length spectra.
The question of whether there exist iso-length-spectral manifolds that are not isometric has a long history. The first examples of such objects are due to Milnor (see [Mil64]), who constructed two flat, 16-dimensional non-isometric, iso-length-spectral tori. In [Vig80], Vignerás constructed examples of hyperbolic surfaces that are iso-length-spectral but not isometric. Both Vignerás’s and Milnor’s constructions are heavily number theoretic and, one could argue, difficult to describe geometrically. In [Sun85], Sunada developed a celebrated, systematic, and more geometric way of constructing hyperbolic iso-length-spectral surfaces. In [Bus86] (and [Bus92] for the genus $6$ case), Buser uses [Sun85] to construct pairs of iso-length-spectral, non-isometric surfaces of genus $\geq 5$.
@@
We remark that in contrast to the surface amalgam case, there cannot exist non-homeomorphic iso-length-spectral pairs of hyperbolic closed surfaces, as the length spectrum completely determines the genus of a surface (see [Cha84]). In the setting of $3$-manifolds, on the other hand, applications of the Sunada method (e.g. [Spa89], [Rei92], [McR14]) have been shown to yield hyperbolic, iso-length-spectral 3-manifolds which are non-isometric and thus non-homeomorphic due to Mostow Rigidity. This provides further evidence towards the fact that surface amalgams share characteristics of both surfaces and 3-manifolds (see also [HST20]).
@@
We follow the terminology from [PR15] and define the weak length spectrum of a locally CAT(1-1) metric space to be a collection of lengths of closed geodesics without multiplicity. If two metric spaces have the same weak length spectrum, we say they are weak length isospectral. Notice that since the weak length spectrum is a subset of the length spectrum, weak length isospectrality is a weaker condition than length isospectrality.
@@
Acknowledgements. I am deeply indebted to Chris Leininger for suggesting the project, discussing it in depth with me, and pointing out gaps in previous drafts. I also thank Alan Reid for discussions that inspired Theorem 1.3, and for pointing out various helpful references. Finally, I thank my doctoral advisors Tullia Dymarz and Caglar Uyanik for helpful discussions and moral support. I also thank the referee for a careful reading of the paper, and for suggestions on improving exposition. This work was partially supported by an NSF RTG grant DMS-2230900.
@@ 2 Buser’s Techniques
We rely heavily on the techniques from [Bus86], which we briefly sketch.
2.1 Buser’s original construction
Buser constructs two iso-length-spectral but non-isometric hyperbolic genus 5 surfaces by gluing together identical copies of right-angled octagons, which he calls building blocks. The gluing scheme is shown below in Figure 1. His construction is modeled off Sunada’s construction using almost conjugate subgroups of $(\mathbb{Z}/8\mathbb{Z})^{\times}\ltimes(\mathbb{Z}/8\mathbb{Z})^{+}$ given by Gerst in [Ger70] (see Example 1 of Section 1 of [Sun85]).
@@ 2.1 Buser’s original construction
Let $0<c<b<1$. Then any geodesic curve $\delta$ of $B_{i}$ which connects two sides of $B_{i}$ has length $\ell(\delta)\geq c$. Equality holds only if $\delta$ is a side (of length $c$) of a building block.
@@
Buser’s surfaces are iso-length-spectral. Next, for every closed geodesic in $S_{1}$, Buser constructs one in $S_{2}$ with the same length (and vice versa) using a technique he calls transplantation. Given a closed geodesic $\gamma\subset S_{1}$ with a starting point $p$ in the interior of some building block $B_{i}\subset S_{1}$, Buser specifies which building block $B^{\prime}_{k(i)}\subset S_{2}$ one should start constructing $\gamma^{\prime}\subset S_{2}$ in so that $\ell(\gamma)=\ell(\gamma^{\prime})$. He uses the following set of rules, which depend on the parities of #a\#a and #b\#b, defined below:
@@
Let #a\#a and #b\#b be the number of times a curve $\gamma$ transversely crosses sides of lengths $a$ and $b$ respectively, and let $B_{n_{1}}$ (resp. $B^{\prime}_{n^{\prime}_{1}}$) denote the building blocks in $S_{1}$ (resp. $S_{2}$) in which to initiate $\gamma$ (resp. $\gamma^{\prime}$).
@@
    If #a\#a is even, take $n_{1}=n^{\prime}_{1}$;
@@
    If #a\#a is odd and #b\#b is even, take $n^{\prime}_{1}=n_{1}+1$;
@@
    If #a\#a and #b\#b are both odd, take $n^{\prime}_{1}=n_{1}+2$.
@@
To construct $\gamma^{\prime}$, one simply decomposes $\gamma$ into geodesic segments $\{\gamma_{j}\}_{j=1}^{N}$ such that each $\gamma_{j}$ is contained completely in some building block $B_{j}$ and $\gamma_{1}$ and $\gamma_{N}$ are contained in the same building block and share an endpoint at $p$. Then, for each $\gamma_{j}$, taking advantage of the fact that the building blocks in $S_{1}$ and $S_{2}$ are all identical, one can construct a $\gamma^{\prime}_{j}\subset B^{\prime}_{j}$ identical to each $\gamma_{j}\subset B_{j}$. Buser shows that following Algorithm 2.2, the set of $\{\gamma^{\prime}_{j}\}$ will always close up to a geodesic loop. Furthermore, since $\ell(\gamma_{j})=\ell(\gamma^{\prime}_{j})$ for all $j$, it follows that $\ell(\gamma)=\ell(\gamma^{\prime})$, as desired. The same set of initiation rules also applies in the other direction (constructing a geodesic in $S_{1}$ given one in $S_{2}$).
2.2 Buser’s techniques in the surface amalgam setting
We now specify some notation and establish some facts used in the iso-length-spectrality proofs in the remainder of the paper. The results and definitions in this section are either heavily inspired by or taken directly from [Bus86].
Let $\beta_{i}$ (resp. $\beta^{\prime}_{i}$) be a connected geodesic segment in $S_{1}$ (resp. $S_{2}$) which can be written as a union $\bigcup\limits_{k=1}^{L}\gamma_{i,k}$ (resp. $\bigcup\limits_{k=1}^{L}\gamma^{\prime}_{i,k}$ ) of segments each completely contained in a single building block. We will see later that every closed geodesic in a surface amalgam constructed in this paper can be written as a concatenation of $\beta_{i}$’s or $\beta^{\prime}_{i}$’s. We define
$\delta_{i}(k)=n^{\prime}_{i,k}-n_{i,k}\pmod{8}$, (1)
where $n_{i,k}$ (resp. $n^{\prime}_{i,k}$) is the index of the building block containing $\gamma_{i,k}$ (resp. $\gamma^{\prime}_{i,k}$). Thus, when we say “initiate $\beta^{\prime}_{i}$ with the rule $\delta_{i}(0)=N$,” we mean that we will set $n^{\prime}_{i,0}$ equal to $n_{i,0}+N$. In other words, if $\beta_{i}$ starts in $B_{n_{i,0}}\subset S_{1}$, then $\beta^{\prime}_{i}$ will start in $B^{\prime}_{n_{i,0}+N}\subset S_{2}$.
For the convenience of the reader, we list observations from Section 5 of [Bus86] used to prove the validity of Algorithm 2.2, which may be checked. We will also use these observations extensively in iso-length-spectrality proofs in the remainder of the paper. Note that while Buser works with closed surfaces, the segments $\beta_{i}$ and $\beta^{\prime}_{i}$ are contained entirely within closed surfaces, so the observations are still applicable in the setting of surface amalgams.
Lemma 2.3 (c.f. Proof of 5.3 in [Bus86]).
Let $\delta_{i}$, $\beta_{i}$, and $\beta^{\prime}_{i}$ be as above. Then the following changes to $\delta_{i}$ are observed whenever $\beta_{i}$ and $\beta^{\prime}_{i}$ cross edges of building blocks of $S_{1}$ and $S_{2}$ respectively:
  1. 1.
    Crossing a side of length $a$. If $\delta_{i}(k)$ is even, then $\delta_{i}(k+1)=\delta_{i}(k)+4\pmod{8}$. If $\delta_{i}(k)$ is odd, then $\delta_{i}(k+1)=\delta_{i}(k)$.
  2. 2.
    Crossing a side of length $b$. If $\delta_{i}(k)=0$ or $4$, $\delta_{i}(k+1)=\delta_{i}(k)$. If $\delta_{i}(k)=\pm 2$, $\delta_{i}(k+1)=\delta_{i}(k)+4\pmod{8}$. There are other possible scenarios, but only these are used in this paper.
  3. 3.
    Crossing a side of length $c$. Regardless of the value of $\delta_{i}(k)$, $\delta_{i}(k+1)=\delta_{i}(k)$.
2.2.1 Translated and transplanted copies
Next, we define translated copies of curves, which, roughly speaking, are locally isometric copies of a curve on the same surface (either $S_{1}$ or $S_{2}$). In contrast, a transplanted copy of a curve in $S_{1}$ is a curve on the other surface $S_{2}$ which has the same length (and vice versa). To formalize this, we first present a definition from [Bus86]:
Definition 2.4 (c.f. Definition 5.2, [Bus86]).
Let $\beta$ and $\beta^{\prime}$ be two curves in $S_{1}$ or $S_{2}$ which have the same length. Suppose $\beta=\bigcup\limits_{k=1}^{L}\beta_{k}$, where each $\beta_{k}$ is a curve contained entirely in a single building block. Similarly, suppose $\beta^{\prime}$ can be written as the union $\beta^{\prime}=\bigcup\limits_{k=1}^{L}\beta^{\prime}_{k}$. Then $\beta$ and $\beta^{\prime}$ are locally congruent if there exist local isometries $\{\varphi_{k}\}_{k=1}^{L}$ such that $\varphi_{k}(N(\beta_{k}))=N(\beta^{\prime}_{k})$ for every $k\in[1,L]$, where $N(\beta_{k})$ and $N(\beta^{\prime}_{k})$ are neighborhoods of $\beta_{k}$ and $\beta^{\prime}_{k}$ respectively.
We can then define translated copies:
Definition 2.5 (Translated copies of geodesic segments).
Given a geodesic segment $\beta=\bigcup\limits_{k=1}^{L}\beta_{k}$ in $S_{1}$, a translated copy of $\beta$ is a curve $\alpha=\bigcup\limits_{k=1}^{L}\alpha_{k}\subset S_{1}$ which is locally congruent to $\beta$. We also require that if $\beta$ begins (resp. ends) on an edge with a certain label, then $\alpha$ begins (resp. ends) on an edge with the same label. Furthermore, if $\beta_{k}$ and $\alpha_{k}$ are contained in the building blocks $B_{n(k)}$ and $B_{m(k)}$ respectively, then for every $1\leq k\leq L-1$, $B_{m(k)}$ meets $B_{m(k+1)}$ along an edge with the same label as that of the edge where $B_{n(k)}$ and $B_{n(k+1)}$ meet. One can also replace $S_{1}$ in the definition with $S_{2}$.
Following terminology from Section 11.6 of [Bus92], we also define the following.
Definition 2.6 (Transplanted copies of geodesic segments).
Given a geodesic segment $\beta=\bigcup\limits_{k=1}^{L}\beta_{k}$ in $S_{1}$, a transplanted copy of $\beta$ is a curve $\beta^{\prime}=\bigcup\limits_{k=1}^{L}\beta^{\prime}_{k}\subset S_{2}$ which is locally congruent to $\beta$. We also require that if $\beta$ begins (resp. ends) on an edge with a certain label, then $\beta^{\prime}$ begins (resp. ends) on an edge with the same label. Furthermore, if $\beta_{k}$ and $\beta^{\prime}_{k}$ are contained in the building blocks $B_{n(k)}\subset S_{1}$ and $B^{\prime}_{n(k)}\subset S_{2}$ respectively, then for every $1\leq k\leq L-1$, $B^{\prime}_{n(k)}$ meets $B^{\prime}_{n(k+1)}$ along an edge with the same label as that of the edge where $B_{n(k)}$ and $B_{n(k+1)}$ meet.
@@ 3.1 Homeomorphic, non-isometric surface amalgams
Consider $S_{1}$ and $S_{2}$ from [Bus86]. In each surface, consider the two closed geodesics of length $2a$ obtained from concatenating the top right edge of length $a$ on $B_{2}$ (resp. $B_{6}$) with the top left edge of length $a$ on $B_{3}$ (resp. $B_{7}$). We do the same for the bottom edges. We then identify all the red edges of length $a$ on $B_{2}$ and $B_{6}$ (resp. $B^{\prime}_{2}$ and $B^{\prime}_{6}$) from Figure 2 and all blue edges on $B_{3}$ and $B_{7}$ (resp. $B^{\prime}_{3}$ and $B^{\prime}_{7}$).
Figure 2: Homeomorphic, non-isometric, iso-length-spectral surface amalgams. The gluing curve consists of two closed geodesics of length $2a$ identified together. Orientations of the gluing curves are specified in the figure. Geodesic segments that are the same colors in $X_{1}$ are identified; the same is true for $X_{2}$. Cutting along the gluing geodesics and the systoles yields two connected components (in yellow and blue) for $X_{1}$ but only one connected component for $X_{2}$.
@@
$X_{1}$ and $X_{2}$ are iso-length-spectral. It suffices to consider geodesics that do not intersect the gluing curve, as [Bus86] already shows a 1-1 correspondence between closed geodesics completely contained in $S_{1}$ and $S_{2}$. Consider $\gamma$, a closed geodesic in $X_{1}$. Decompose $\gamma$ into $\bigcup\limits_{i=1}^{N}\beta_{i}$ so that each $\beta_{i}$ is a continuous geodesic segment in the original surface $S_{1}$ with endpoints on the gluing curve. We say a closed geodesic in $\gamma_{1}\subset X_{1}$ is identical to $\gamma$ if $\gamma_{1}$ and $\gamma$ begin at the same point $x\in X_{1}$ on the gluing curve, and $\gamma_{1}$ can be decomposed into $\bigcup\limits_{i=1}^{N}\beta_{i}^{1}$, a union of translated copies of $\beta_{i}$ in $S_{1}$ with endpoints on the gluing curve. There is a natural isometry $\varphi$ between gluing geodesics in $X_{1}$ and those in $X_{2}$. We say $\gamma^{\prime}\subset X_{2}$ is identical to $\gamma\subset X_{1}$ if $\gamma^{\prime}$ begins and ends at $\varphi(x)\in X_{2}$ while $\gamma$ begins and ends at $x\in X_{1}$, and $\gamma_{2}$ can be decomposed into $\bigcup\limits_{i=1}^{N}\beta^{\prime}_{i}$ so that each $\beta^{\prime}_{i}$ is a transplanted copy of $\beta_{i}$ (in the sense of [Bus86]) that begins and ends on the gluing curve in $X_{2}$. We will argue that the numbers of closed geodesics identical to $\gamma$ in $X_{1}$ and $X_{2}$ are the same, which shows the length spectra are the same.
@@
Figure 4: Two examples of (parts of) inadmissible $\beta_{i}$’s. In (a), $\beta_{i-1}\cup\beta_{i}$ is a backtracking geodesic segment. In (b), $\beta_{i-1}$ and $\beta_{i}$ begin at different points of the gluing curve, as the red and blue segments are not identified, which means $\beta_{i-1}\cup\beta_{i}$ is disconnected.
We first prove the following:
Claim. In $X_{1}$, there is exactly one translate of $\beta_{i}^{1}:=\beta_{i}$, $\beta_{i}^{2}$, which begins and ends at the same point on the gluing curve $\beta_{i}$ begins and ends at.
Proof of claim: If $\beta_{i}$ begins on the top right (resp. top left, bottom right, or bottom left) side of length $a$ in $B_{i}$, initiate $\beta_{i}^{2}$ at the same point on the top right (resp. top left, bottom right, or bottom left) side of length $a$ in $B_{i+4}$. Let $\beta_{i}^{j}=\bigcup\limits_{k=0}^{L}\gamma_{i,k}^{j}$ be a decomposition of $\beta_{i}^{j}$ into geodesic segments that are contained entirely in a single building block, where $\gamma_{i,k}^{j}\subset B_{n_{i,k}^{j}}$ and $j=1,2$. Let $\epsilon_{k}=n_{i,k}^{2}-n_{i,k}^{1}.$ Note in particular that $\epsilon_{0}=4$. We observe the following changes to $\epsilon_{j}$ in the following scenarios:
  • $\beta^{1}_{i}$ and $\beta^{2}_{i}$ cross sides of length $c$. In this case, $n^{j}_{i,k+1}=n^{j}_{i,k}\pm 1$ for both $j=1,2$ depending on whether the left or right side of a building block is crossed. Note that when crossing edges of length $c$, $\beta^{1}_{i}$ and $\beta^{2}_{i}$ either both cross left sides or both cross right sides. Then $\epsilon_{k+1}=(n^{2}_{i,k}\pm 1)-(n^{1}_{i,k}\pm 1)=n^{2}_{i,k}-n^{1}_{i,k}=\epsilon_{k}$.
  • $\beta^{1}_{i}$ and $\beta^{2}_{i}$ cross sides of length $a$. In this case, for both $j=1,2$, $n^{j}_{i,k+1}=n^{j}_{i,k}$ or $n^{j}_{i,k}+4$, depending on the parity of the building block. Provided that $\epsilon_{k}=4\pmod{8}$, either $\epsilon_{k+1}=n^{2}_{i,k}-n^{1}_{i,k}=\epsilon_{k}$ or $\epsilon_{k+1}=(n^{2}_{i,k}+4)-(n^{1}_{i,k}+4)=\epsilon_{k}.$
  • $\beta^{1}_{i}$ and $\beta^{2}_{i}$ cross sides of length $b$. In this case, for both $j=1,2$, $n^{j}_{i,k+1}=n^{j}_{i,k}\pm 1$ or $n^{j}_{i,k}\pm 3$, depending on the label of the building block. But again, if $\epsilon_{k}=4$, then $\epsilon_{k+1}=(n^{2}_{i,k}\pm 1)-(n^{1}_{i,k}\pm 1)=n^{2}_{i,k}-n^{1}_{i,k}=\epsilon_{k}$ or $\epsilon_{k+1}=(n^{2}_{i,k}\pm 3)-(n^{1}_{i,k}\pm 3)=n^{2}_{i,k}-n^{1}_{i,k}=\epsilon_{k}$.
From this, we conclude that in fact, $\epsilon_{L}=4$. That is, if $\beta^{1}_{i}$ ends on an edge of the building block $B_{N}$, then $\beta^{2}_{i}$ ends on the same point in the corresponding edge of $B_{N+4}$. Since there are only two edges that are top right, top left, bottom right, or bottom left edges of building blocks which are identified to create a subarc of the gluing geodesic, there cannot be another translate of $\beta^{1}_{i}$ sharing beginning and end points with $\beta^{1}_{i}$. Thus, there is exactly one such $\beta^{2}_{i}$, as claimed. ∎
@@
[…]
We now examine how to choose a copy of $\beta_{N}$ from $\{\beta_{N}^{j}\}_{j=1}^{2}$. The set of admissible $\beta_{N}$ now depends on our choice of $\beta_{N-1}^{j}$ and $\beta_{1}^{1}$, and specifically on whether the following facts are true:
  • Fact 1: $\beta^{j}_{N-1}$ ends on a top red edge and $\beta_{N}$ begins on a top red edge (or the statement is true if we replace “red” with “blue” and/or “top” with “bottom”);
@@
Case One: Neither Fact 1 nor Fact 2 is true. In this case, any choice of pairs of translates of $\beta_{1}$ and $\beta_{N}$ is compatible. Thus, each choice of $\beta_{1}^{j}$ ($j=1,2$) has $\prod\limits_{i=2}^{N}c_{i}=\bigg{(}\prod\limits_{i=2}^{N-1}c_{i}\bigg{)}(2)$ choices of sequences of admissible translates of $\beta_{i}$. Then $C_{1}(\gamma)=2\bigg{(}\prod\limits_{i=2}^{N-1}c_{i}\bigg{)}(2)$. See Figure 5.
Figure 5: An illustration of Case 1, with identical copies of closed geodesics created by concatenating the pink ($\beta_{1})$, green ($\beta_{2}$), and orange ($\beta_{3}$) geodesic segments. Two closed geodesics are illustrated in each of $X_{1}$ and $X_{2}$, one solid and one dashed. The star icon indicates the start of each geodesic. Here, $N=3$ and $c_{2}=2$, so $C_{1}(\gamma)=2(2)(2)=C_{2}(\gamma)$. We can check that there are indeed eight identical, non-backtracking closed geodesics in each of $X_{1}$ and $X_{2}$.
@@
Figure 6: An illustration for Case 2. Again, two closed geodesics are shown, one solid and one dashed, and the star indicates the starting point of each geodesic. Here, $N=3$ and $c_{2}=2$, so $C_{1}(\gamma)=2(2)=C_{2}(\gamma)$. We can check that there are four identical, non-backtracking closed geodesics in each of $X_{1}$ and $X_{2}$.